1. A baseball runner is running west towards 2nd base at 7.2 m/s and 0.80 s later has rounded 2nd and is heading north towards 3rd base at 8.4 m/s. Determine his change in velocity, and his acceleration.
1
Expert's answer
2021-05-06T07:41:50-0400
"\Delta v=\sqrt{7.2^2+8.4^2}=11.1\frac{m}{s}\\\theta=\arctan{\frac{8.4}{7.2}}=49.4\degree \text{ N of W}\\a=\frac{\sqrt{7.2^2+8.4^2}}{0.8}=13.8\frac{m}{s^2}"